OpenStudy (anonymous):

Given the linear equation y=-7/4x-1, find the y-coordinates of the points (-4, )(0, ) and )4, )... Please show all work... YOU WILL GET THE MEDAL.

OpenStudy (anonymous):

You can find the y coordinate by substituting in the x coordinate given into the equation. Here I'll show you the first one. \[y =( -7/ 4)(-4) -1\] y= 7 -1 y =6 So the first pair would be (-4,6) Now try the next two :)

OpenStudy (anonymous):

so the 4's cancels eachother out

OpenStudy (aroub):

yes! :)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now, try when x = 0 for the second coordinate point.

OpenStudy (anonymous):

y=(-7/4)(0)-1 y=0-1 y=1 right

OpenStudy (anonymous):

almost, very close. the negative doesn't disappear. so y = -1 giving you a coordinate point of (0,-1)

OpenStudy (anonymous):

okay,and the same for the last one as done for the first one but with a negative. right

OpenStudy (anonymous):

Yes after substituting the 4. You are left with y = -7 - 1

OpenStudy (anonymous):

ok thank you for teaching me you get the medal